# Gradient Diffusion: A Perturbation-Resilient Gradient Leakage Attack

### v1

@@ 3 Methodology
$q(X_{t-1}|X_{0},X_{t})=N(X_{t};\mu_{t},\sigma_{t}I).$ (12)
Algorithm 3 Generate Original Gradient
1:  if Known Noise Scale $M$ then
2:     $\textbf{c}=\frac{1}{\sqrt{1+M^{2}}}$;
3:  else
4:     $\quad\textbf{c}\in(0,1)$;
5:  end if
6:  $X_{T^{\prime}}\sim\textbf{c}\nabla W^{\prime}$;
7:  for $t=T^{\prime},...,1$ do
8:     if $t>1$ then
9:        $\textbf{z}\sim N(\textbf{0},\textbf{I})$;
10:     else
11:        $\quad\textbf{z}=\textbf{0}$;
12:     end if
13:     $X_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(X_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}f_{\theta}(X_{t},t))+\sqrt{1-\alpha_{t}}\textbf{z}$;
14:  end for
15:  Return $X_{0}$;
16:  $\nabla W^{R}\leftarrow X_{0}$; (For Further Gradient-Based Attack)
If insert the stolen Shared Perturbed Gradients $\nabla W^{\prime}$ or constructed Surrogate Perturbed Gradients $\nabla W^{s}_{perturbed}$ as training condition then input $X_{0}=(condition,\nabla W_{j})$.If not to train with the condition, then input $X_{0}=\nabla W_{j}$.Through algebraic calculation, $\mu_{t}$ and $\sigma_{t}$ in Eq. (12) can be simplified for further usage in the reverse training process[24] as:
$\mu_{t}$ $=$ $\frac{\sqrt{\gamma_{t-1}}(1-\alpha_{t})}{1-\gamma_{t}}X_{0}+\frac{\alpha_{t}(1-\gamma_{t-1})}{1-\gamma_{t}}X_{t},$ (13)
$\sigma^{2}_{t}$ $=$ $\frac{1-\gamma_{t-1}}{1-\gamma_{t}}(1-\alpha_{t}).$ (14)
In the reverse process $(p)$ of PGLA’s Gradient Diffusion training, the objective function $f_{\theta}$ which is trained to predict noise vector $\epsilon$ is modeled as:
$E_{X_{0},\epsilon}[\frac{(1-\alpha_{t})^{2}}{2\sigma^{2}_{t}\alpha_{t}(1-\gamma_{t})}||f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)-\epsilon||^{2}].$ (15)
In Step (4), refer to Algorithm 3 lines 1 to 4, two different original gradient generation processes are chosen depending on whether or not attackers know the Noise Scale of Perturbation of the Threat Model. Take FL with Gaussian Differential Privacy (Eq. (6) in Preliminary) as an example, consider an experienced and clever attacker who may know the DP Privacy Budget $\varepsilon$ and the probability of information leakage $\delta$ ($\delta$ is likely to be set as $10^{-5}$ from usual practice in DP), combining with the sensitivity $\nabla S$ (which can be estimated if the attacker has some previous information with the target model training dataset), the noise scale can be calculated or estimated to an approximate value $M$ ($M$ defined as the adaptive parameter in [M-Adaptive Process]). So, $\nabla W^{\prime}$ can be modeled as:
$\nabla W^{\prime}=\nabla W+MN(0,I).$ (16)
@@
$\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}=\frac{1}{\sqrt{1+M^{2}}}\nabla W+\frac{M}{\sqrt{1+M^{2}}}N(0,I).$ (17)
Recall that our target is to generate $\nabla W^{R}$, so $\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}$ should be considered as $X_{t}$ in the inverse process of Gradient Diffusion Model, while $\nabla W$ stands for $X_{0}$. Correspondingly, during the construction of the recovered gradient $\nabla W^{R}$, the inverse time steps can be set as $T^{\prime}$. The relationship between $T^{\prime}$ and $M$ is:
$\frac{1}{1+M^{2}}=\prod_{t=0}^{T^{\prime}}\alpha_{t}.$ (18)
@@
4 Experiments
We assess the performance of the PGLA within the FL-PP framework, which encompasses the FL-DP approach.
4.1 Experimental Setups.
@@ 4 Experiments
Variant Attack Models Three PGLA variant models: PGLA (trained with only unperturbed surrogate gradients), Condition PGLA (trained with both perturbed gradients and unperturbed surrogate gradients), and Non-Adaptive PGLA (without [M-Adaptive Process]: no perturbation scale $M$ as adaptive parameter during gradient diffusion process) are presented during experiments.
@@
Datasets In the subsequent experimental analyses, we employ the datasets MNIST, CIFAR100, and STL10 as client privacy datasets, which also serve as the ground truth for privacy leakage evaluation. We extract the unperturbed original gradients ($\nabla W$) of the aforementioned three datasets from the local training model of the target client as the reference benchmark of gradient denoising under the FL-PP paradigm.
Evaluation and Boundary In the context of evaluating Gradient Denoising (Sec. 4.2 $\sim$ 4.3), we employ the cosine similarity $CosSimilar_{g}$ and the Average Peak Signal-to-Noise Ratio $PSNR_{g}$ as metrics to assess the quality of the Recovered Gradients $\nabla W^{R}$ compared to the Original Gradients $\nabla W$. Higher values of the two metrics indicate better accuracy in the original gradient recovery.
@@
To evaluate the Privacy Leakage Capability (Sec. 4.4), we utilize the Image Average Peak Signal-to-Noise Ratio $PSNR_{i}$ and the Label Recovered Accuracy $LRA$. These metrics are employed to assess the fidelity of the recovered images and the accuracy of the recovered labels, respectively, to the original images and ground truth labels. The boundary of Privacy Leakage Attack is aligned with previous works (Sec. 2.4): the attack is considered successful if human visual perception can discern the requisite information from the recovered images.
4.2 Gradients Denoising under FL-DP
Among FL-DP, we choose the NbAFL framework [28] (Noising before model aggregation FL, details refer to Sec. 2.2) as the threat model in this experiment due to its widespread adoption. To ensure a comprehensive evaluation of the effectiveness of the PGLA, we train the three PGLA variant models, as well as non-diffusion denoising models such as NBNet [3], SS-BSN [11], and AP-BSN [14], using gradients extracted from the surrogate downstream task model trained on the FashionMnist. Following that, we utilize the trained models to denoise the shared perturbed gradients intercepted by the attacker from the target clients. These target clients have locally trained downstream task models using privacy datasets (MNIST, CIFAR100, and STL10). The gradients are protected using the NbAFL[28] before being shared with the server.
Figure 4: The visualization of gradients denoising on the MNIST dataset.
Based on the experimental results presented in Table 2, it is evident that PGLA showcases superior denoising ability for perturbed gradients under FL-DP, with an average Cosine Similarity exceeding 0.992 and a PSNR of 37.68 on the three datasets. This represents a significant improvement of over $27\%$ compared to non-diffusion methods. Fig.4 visualizes the denoising process of gradients with different methods, we can also observe that PGLA has a better denoising effect compared to NBNET, while SS-BSN and AP-BSN destroy the original structure of the gradient due to their over-denoise effect.
Further analysis reveals that among the three PGLA variant models, the performance can be ranked as follows: PGLA(Condition)$>$ PGLA $>$ PGLA (Non-Adaptive). This ranking demonstrates both M-Adaptive Process and conditional training enhance the denoising and generation performance of PGLA’s Gradient Diffusion Model. It suggests that selecting an appropriate PGLA variant model based on available prior information can improve attack performance.
@@ 4.2 Gradients Denoising under FL-DP
Noised Gradient 0.992 35.74 0.979 27.51 0.979 27.23
NBNet[3] 0.792 18.07 0.764 15.48 0.768 15.61
SS-BSN[11] 0.971 29.82 0.845 22.85 0.845 22.84
AP-BSN[14] 0.968 29.61 0.892 24.01 0.893 24.12
PGLA 0.996 38.76 0.990 30.42 0.990 30.01
PGLA(NA) 0.995 38.59 0.990 30.39 0.990 29.87
PGLA(C) 0.996 38.78 0.990 30.53 0.993 30.22
4.3 Gradients Denoising under FL-PP
To evaluate the generalization capability of PGLA on gradient denoising, we constructed two different types of noise (Laplace and Gaussian) randomly applied to each layer of the original gradients within the FL-PP framework. The training and inference processes of PGLA variant models, and non-diffusion denoising models (NBNet [3], SS-BSN [11], and AP-BSN [14]) are the same as the above experiments on gradients denoising under FL-DP. In contrast to attacks specifically tailored for the FL-DP framework, the gradient denoising experiments on FL-PP showcase PGLA’s ability to effectively handle various types of perturbations and adapt to different magnitudes of perturbation.
Referring to the experimental results illustrated in Fig. 5, it is observed that when subjected to Laplace perturbation, PGLA variant models exhibit an average improvement of $21.6\%$ in PSNR and $9.2\%$ in Cosine Similarity. Similarly, under Gaussian perturbation, PGLA achieves an average enhancement of $21.3\%$ in PSNR and $9.1\%$ in Cosine Similarity. These findings provide compelling evidence for the robustness and stability of the PGLA and its variant models.
@@ 4.3 Gradients Denoising under FL-PP
Figure 5: Gradients denoising under FL-PP: Gaussian and Laplace perturbed gradients denoising performance of $PSNR_{g}$ and $CosSimilar_{g}$ via different perturbation magnitudes. The same noise magnitude parameter $\varepsilon$ as in FL-DP is used to represent different levels of perturbation. Perturbation decreases, $\varepsilon$ increases. (Private dataset: STL10)
4.4 Privacy Leakage Capability
The comparison of overall privacy leakage capability from perturbed gradients of PGLA and typical Gradient Leakage Attacks (GRNN [22], IG [9], and DLG [35]) under FL-DP, compares Image Average Peak Signal Noise Ratio $PSNR_{i}$ and Label Recovered Accuracy $LRA$ of the recovered local clients’ privacy information. Local clients’ private training datasets are MNIST, CIFAR100, and STL10.
Table 3: Privacy leakage capability of PGLA variant models and traditional Gradients Leakage Attacks under FL-DP ($\delta$ = $10^{-5}$, $\varepsilon$ = 1, 5, 10). NA: Non-Adaptive; C: Condition; Gray marker: PGLA outperforms the highest result of traditional ones.
$PSNR_{i}$ LRA $PSNR_{i}$ LRA $PSNR_{i}$ LRA
MINIST GRNN[22] 14.70 1.000 16.54 1.000 20.19 1.000
IG[9] 3.721 - 5.164 - 6.067 -
DLG[35] 0.222 0.676 8.025 0.752 17.08 0.909
PGLA 17.87 0.851 24.17 0.895 33.09 0.927
PGLA(NA) 17.39 0.851 24.14 0.889 32.98 0.925
PGLA(C) 20.87 0.887 28.39 0.924 35.14 0.957
CIFAR100 GRNN[22] 18.24 1.000 20.61 1.000 21.15 1.000
IG[9] 3.684 - 5.073 - 5.971 -
DLG[35] 0.112 0.769 7.673 0.833 18.57 0.896
PGLA 18.25 0.883 19.04 0.902 21.32 0.902
PGLA(NA) 18.15 0.871 17.98 0.884 21.06 0.891
PGLA(C) 18.69 0.901 21.02 0.902 23.49 0.908
STL10 GRNN[22] 12.24 1.000 16.24 1.000 20.98 1.000
IG[9] 4.323 - 6.082 - 7.851 -
DLG[35] 0.038 0.783 5.673 0.815 17.25 0.851
PGLA 12.36 0.806 16.30 0.857 19.77 0.889
PGLA(NA) 12.28 0.801 15.19 0.849 19.76 0.887
PGLA(C) 12.98 0.816 20.30 0.872 22.77 0.906
Figure 6: Visualization results on data recovery. Comparisons on ground truth images and corresponding recovered privacy images from PGLA variant models and commonly used traditional gradient leakage attacks. ($\delta$ = $10^{-5}$; $\varepsilon=10$)
According to the numerical experimental results shown in Table 3 and the visualization results shown in Fig. 6, PGLA variant models exhibit a substantial superiority over traditional gradient leakage attacks in terms of private image leakage. On average, there is an approximately $209\%$ increase in the recovered image $PSNR_{i}$ when using PGLA. It’s worth noting that GRNN can always accurately recover the label, keeping its LRA metric at 1.000. The reason is that GRNN has uniquely designed a robust generative network specifically to recover labels, while other methods obtain the label by updating the image and label jointly with a single model during the backward phase, leading to joint image and label recovery. Moreover, upon examining the private image recovery procedures of PGLA variant models and traditional gradient leakage attacks illustrated in Fig. 7, it becomes evident that the attacks incorporating PGLA not only achieve considerably improved accuracy in the ultimate reconstruction of private images compared to conventional approaches but also showcase notable advantages in terms of attack iteration rounds and speed (PGLA (condition) $>$ PGLA $>$ PGLA (Non-Adaptive)). Above all, PGLA variant models demonstrate significant advantages in recovering private images.
Figure 7: Comparison of the private image recovery procedures to the iterations between PGLA variant models and traditional Gradient Leakage Attack methods (DLG[35]).
4.5 Attack Duration
To comprehensively evaluate the attack efficiency of PGLA, we compared the average inference time of different denoising models under FL-PP. The training and inference procedures of the variant models of PGLA, as well as the non-diffusion denoising models, are configured to be consistent as in Sec. 4.3.
The experimental results presented in Table 4 indicate that PGLA variant models exhibit a relative advantage in terms of gradient denoising speed, surpassing non-diffusion methods by an average improvement of $32.7\%$ in inference time.
Limitations: The experiments demonstrate that PGLA can reconstruct original gradients from noise-perturbed gradients without requiring knowledge of the type and magnitude of perturbation noise, as well as FL clients’ local model. However, PGLA is not effective in attacking perturbations that are not based on gradient diffusion (noise perturbation) such as representation perturbation. In addition, the overall effectiveness of PGLA in reconstructing privacy datasets is also bounded by the selected subsequent Gradient Leakage Attacks.
@@ 4.5 Attack Duration
Table 4: Gradient denoising average inference time of PGLA variant models and non-diffusion denoising models under FL-PP. (Private dataset: STL10, $\varepsilon$ = 0.5 $\sim$ 10)
Model NBNet [3] SS-BSN [11] AP-BSN[14] PGLA PGLA(NA) PGLA(C)
Times (s) 2.653 26.15 7.138 6.834 6.148 6.917
@@
Defense Strategies: Possible defense strategies against PGLA can be approached from preserving the privacy of original data, preserving the target attack models, and shared gradients protection. PGLA effectively leakages privacy data by attacking the perturbed shared gradients, which suggests the possible defense approaches against PGLA should either be non-perturbation gradient protection methods or non-gradient privacy-preserving methods. For instance, processing of the participant’s local training model in depth and layers to weaken the performance of Gradient Leakage Attacks, encryption on original local privacy data before training, and various cryptography-based protections on shared gradients [13, 32, 23]. Cryptography-based approaches always cause more computation time and storage [31], thus lightweight defense strategies that can serve as substitutes for invalid perturbation protection still need future explorations.
5 Conclusion

### v2

@@ Introduction
$q(X_{t-1}|X_{0},X_{t})=N(X_{t};\mu_{t},\sigma_{t}I)$ (10)
If insert the stolen Shared Perturbed Gradients $\nabla W^{\prime}$ or constructed Surrogate Perturbed Gradients $\nabla W^{s}_{perturbed}$ as training condition then input $X_{0}=(condition,\nabla W_{j})$.If not to train with the condition, then input $X_{0}=\nabla W_{j}$.Through algebraic calculation, $\mu_{t}$ and $\sigma_{t}$ in Eq. (10) can be simplified for further usage in the reverse training process(Saharia et al. 2023) as:
$\mu_{t}$ $=$ $\frac{\sqrt{\gamma_{t-1}}(1-\alpha_{t})}{1-\gamma_{t}}X_{0}+\frac{\alpha_{t}(1-\gamma_{t-1})}{1-\gamma_{t}}X_{t}$ (11)
$\sigma^{2}_{t}$ $=$ $\frac{1-\gamma_{t-1}}{1-\gamma_{t}}(1-\alpha_{t})$ (12)
In the reverse process $(p)$ of Mjölnir’s Gradient Diffusion training, the objective function $f_{\theta}$ which is trained to predict noise vector $\epsilon$ is modeled as:
$E_{X_{0},\epsilon}[\frac{(1-\alpha_{t})^{2}}{2\sigma^{2}_{t}\alpha_{t}(1-\gamma_{t})}||f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)-\epsilon||^{2}]$ (13)
$PSNR_{i}$ $LRA$ $PSNR_{i}$ $LRA$ $PSNR_{i}$ $LRA$
MINIST GRNN(Ren, Deng, and Xie 2022) 14.70 1.000 16.54 1.000 20.19 1.000
IG(Geiping et al. 2020) 3.721 - 5.164 - 6.067 -
DLG(Zhu, Liu, and Han 2019) 0.222 0.676 8.025 0.752 17.08 0.909
Mjölnir 17.87 0.851 24.17 0.895 33.09 0.927
Mjölnir(Non-Adaptive) 17.39 0.851 24.14 0.889 32.98 0.925
Mjölnir(Condition) 20.87 0.887 28.39 0.924 35.14 0.957
CIFAR100 GRNN(Ren, Deng, and Xie 2022) 18.24 1.000 20.61 1.000 21.15 1.000
IG(Geiping et al. 2020) 3.684 - 5.073 - 5.971 -
DLG(Zhu, Liu, and Han 2019) 0.112 0.769 7.673 0.833 18.57 0.896
Mjölnir 18.25 0.883 19.04 0.902 21.32 0.902
Mjölnir(Non-Adaptive) 18.15 0.871 17.98 0.884 21.06 0.891
Mjölnir(Condition) 18.69 0.901 21.02 0.902 23.49 0.908
STL10 GRNN(Ren, Deng, and Xie 2022) 12.24 1.000 16.24 1.000 20.98 1.000
IG(Geiping et al. 2020) 4.323 - 6.082 - 7.851 -
DLG(Zhu, Liu, and Han 2019) 0.038 0.783 5.673 0.815 17.25 0.851
Mjölnir 12.36 0.806 16.30 0.857 19.77 0.889
Mjölnir(Non-Adaptive) 12.28 0.801 15.19 0.849 19.76 0.887
Mjölnir(Condition) 12.98 0.816 20.30 0.872 22.77 0.906
Table 1: Privacy leakage capability of Mjölnir variant models and traditional Gradients Leakage Attacks under FL-DP ($\delta$ = $10^{-5}$, $\varepsilon$ = 1, 5, 10). NA: Non-Adaptive; C: Condition; Gray marker: Mjölnir outperforms the highest result of traditional ones.
In Step (4) Recover Original Gradients, refer to Algorithm 3, two different original gradient generation processes are chosen depending on whether or not attackers know the Noise Scale of Perturbation of the Threat Model. Take FL with Gaussian Differential Privacy (Eq. (6)) as an example, consider an experienced and clever attacker who may know the DP Privacy Budget $\varepsilon$ and the probability of information leakage $\delta$ ($\delta$ is likely to be set as $10^{-5}$ from usual practice in DP), combining with the sensitivity $\nabla S$ (which can be estimated if the attacker has some previous information with the target model training dataset), the noise scale can be calculated or estimated to an approximate value $M$ ($M$ defined as the adaptive parameter in [M-Adaptive Process]). So, $\nabla W^{\prime}$ can be modeled as:
$\nabla W^{\prime}=\nabla W+MN(0,I)$ (14)
@@
$\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}=\frac{1}{\sqrt{1+M^{2}}}\nabla W+\frac{M}{\sqrt{1+M^{2}}}N(0,I)$ (15)
Recall that our target is to generate $\nabla W^{R}$, so $\frac{1}{\sqrt{1+M^{2}}}\nabla W^{\prime}$ should be considered as $X_{t}$ in the inverse process of Gradient Diffusion Model, while $\nabla W$ stands for $X_{0}$.
Correspondingly, during the construction of the recovered gradient $\nabla W^{R}$, the inverse time steps can be set as $T^{\prime}$. The relationship between $T^{\prime}$ and $M$ is:
$\frac{1}{1+M^{2}}=\prod_{t=0}^{T^{\prime}}\alpha_{t}$ (16)
@@
Figure 4: Comparison of the private image recovery procedures to the iterations between Mjölnir variant models and traditional Gradient Leakage Attack methods (DLG(Zhu, Liu, and Han 2019)).
Figure 5: Comparisons on ground truth clients’ private images and corresponding recovered privacy images from Mjölnir variant models and commonly used traditional gradient leakage attacks. ($\delta$ = $10^{-5}$; $\varepsilon=10$)
Experiments
Experimental Setups
@@
(A) Mjölnir Variant Attack Models. Mjölnir (trained with only unperturbed surrogate gradients), Condition Mjölnir (trained with both perturbed gradients and unperturbed surrogate gradients), and Non-Adaptive Mjölnir (without [M-Adaptive Process]: no perturbation scale $M$ as an adaptive parameter during gradient diffusion process) are presented.
@@
(B) Benchmarks and Datasets. We employ MNIST, CIFAR100, and STL10 as client privacy datasets, which also serve as the ground truth for privacy leakage evaluation. We extract the unperturbed original gradients ($\nabla W$) of the aforementioned three datasets from the local training model of the target client as the reference benchmark of gradient denoising under the FL-PP paradigm. The Mjölnir gradient diffusion model is trained with gradients extracted from a separate dataset, FashionMNIST.
@@
(C) Evaluation and Boundary. To evaluate the Privacy Leakage Capability, we utilize the Image Average Peak Signal-to-Noise Ratio $PSNR_{i}$ and the Label Recovered Accuracy $LRA$. These metrics are employed to assess the fidelity of the recovered images and the accuracy of the recovered labels, respectively, to the original images and ground truth labels. The boundary of Privacy Leakage Attack is aligned with previous works: the attack is considered successful if human visual perception can discern the requisite information from the recovered images. In the context of evaluating Gradient Denoising, we employ the cosine similarity $CosSimilar_{g}$ and the Average Peak Signal-to-Noise Ratio $PSNR_{g}$ as metrics to assess the quality of the Recovered Gradients $\nabla W^{R}$ compared to the Original Gradients $\nabla W$. Higher values of the two metrics indicate better accuracy in the original gradient recovery.
Privacy Leakage Capability
The comparison of overall privacy leakage capability from perturbed gradients of Mjölnir and traditional Gradient Leakage Attacks (GRNN (Ren, Deng, and Xie 2022), IG (Geiping et al. 2020), and DLG (Zhu, Liu, and Han 2019)) under FL-DP, compares Image Average Peak Signal Noise Ratio $PSNR_{i}$ and Label Recovered Accuracy $LRA$ of the recovered local clients’ privacy information. Local clients’ private training datasets are MNIST, CIFAR100, and STL10.
According to the numerical experimental results shown in Table 1 and the visualization results shown in Fig. 5, Mjölnir variant models exhibit a substantial superiority over traditional gradient leakage attacks in terms of private image leakage. On average, there is an approximately $209\%$ increase in the recovered image $PSNR_{i}$ when using Mjölnir. For traditional methods, GRNN consistently achieves an LRA metric of 1.000 by using a robust generative network for label recovery, while other methods update images and labels jointly, leading to less accurate results. Moreover, upon examining the private image recovery procedures of Mjölnir variant models and traditional gradient leakage attacks illustrated in Fig. 4, it becomes evident that the attacks incorporating Mjölnir not only achieve considerably improved accuracy in the ultimate reconstruction of private images compared to conventional approaches but also showcase notable advantages in terms of attack iteration rounds and speed (Mjölnir (Condition) $>$ Mjölnir $>$ Mjölnir (Non-Adaptive)). Above all, Mjölnir variant models demonstrate significant advantages in recovering private images.
@@
NBNet(Cheng et al. 2021) 0.992 35.74 0.979 27.51 0.979 27.23
SS-BSN(Han and Yu 2023) 0.995 32.35 0.845 22.85 0.845 22.84
AP-BSN(Lee, Son, and Lee 2022) 0.968 29.61 0.892 24.01 0.893 24.12
Mjölnir 0.996 38.76 0.990 30.42 0.990 30.01
Mjölnir (Non-Adaptive) 0.995 38.59 0.990 30.39 0.990 29.87
Mjölnir(Condition) 0.996 38.78 0.990 30.53 0.993 30.22
@@
Figure 6: Gradients denoising under FL-PP: Gaussian and Laplace perturbed gradients denoising performance of $PSNR_{g}$ and $CosSimilar_{g}$ via different perturbation magnitudes. The same noise magnitude parameter $\varepsilon$ as in FL-DP is used to represent different levels of perturbation. Perturbation decreases, $\varepsilon$ increases. (Private dataset: STL10, $\delta$ = $10^{-5}$)
Gradients Denoising under FL-DP
Among FL-DP, we choose the NbAFL framework (Wei et al. 2020) (Noising before model aggregation FL) as the threat model in this experiment due to its widespread adoption. To ensure a comprehensive evaluation of the effectiveness of the Mjölnir, we train the three Mjölnir variant models, as well as non-diffusion denoising models such as NBNet (Cheng et al. 2021), SS-BSN (Han and Yu 2023), and AP-BSN (Lee, Son, and Lee 2022), using gradients extracted from the surrogate downstream task model trained on the FashionMnist. Following that, we utilize the trained models to denoise the shared perturbed gradients intercepted by the attacker from the target clients. These target clients have locally trained downstream task models using privacy datasets (MNIST, CIFAR100, and STL10). The gradients are protected using the NbAFL(Wei et al. 2020) before being shared with the server.
Based on the experimental results presented in Table 2, it is evident that Mjölnir showcases superior denoising ability for perturbed gradients under FL-DP, with an average Cosine Similarity exceeding 0.992 and a PSNR of 37.68. This represents a significant improvement of over $27\%$ compared to non-diffusion methods. Further analysis reveals that among the three Mjölnir variant models, the performance can be ranked as follows: Mjölnir(Condition)$>$ Mjölnir $>$ Mjölnir (Non-Adaptive). This ranking demonstrates both M-Adaptive Process and conditional training enhance the denoising and generation performance of Mjölnir’s Gradient Diffusion Model. It suggests that selecting an appropriate Mjölnir variant model based on available prior information can improve attack performance.
Model Inference time (s)
NBNet(Cheng et al. 2021) 2.653
SS-BSN(Han and Yu 2023) 26.15
AP-BSN(Lee, Son, and Lee 2022) 7.138
Mjölnir 6.834
Mjölnir (Non-Adaptive) 6.834
Mjölnir(Condition) 6.917
@@
Table 3: Gradient denoising average inference time of Mjölnir variant models and non-diffusion denoising models under FL-PP. (Private dataset: STL10, $\varepsilon$ = 0.5 $\sim$ 10, $\delta$ = $10^{-5}$. Device: NVIDIA GeForce RTX 2060 GPU; Intel(R) Core(TM) i7-10870H CPU at 2.20GHz.)
Gradients Denoising under FL-PP
To evaluate the generalization capability of Mjölnir on gradient denoising, we constructed two different types of noise (Laplace and Gaussian) randomly applied to each layer of the original gradients within the FL-PP framework. The training and inference processes of Mjölnir variant models, and non-diffusion denoising models (NBNet (Cheng et al. 2021), SS-BSN (Han and Yu 2023), and AP-BSN (Lee, Son, and Lee 2022)) are the same as the above experiments on gradients denoising under FL-DP. In contrast to attacks specifically tailored for the FL-DP framework, the gradient denoising experiments on FL-PP showcase Mjölnir’s ability to effectively handle various types of perturbations and adapt to different magnitudes of perturbation.
Referring to the experimental results illustrated in Fig. 6, it is observed that when subjected to Laplace perturbation, Mjölnir variant models exhibit an average improvement of $21.6\%$ in PSNR and $9.2\%$ in Cosine Similarity. Similarly, under Gaussian perturbation, Mjölnir achieves an average enhancement of $21.3\%$ in PSNR and $9.1\%$ in Cosine Similarity. These findings provide compelling evidence for the robustness and stability of the Mjölnir variant models.
We further compared the average inference time of different denoising models under FL-PP. The experimental results presented in Table 3 indicate that Mjölnir variant models exhibit a relative advantage in terms of gradient denoising speed, surpassing non-diffusion methods by an average improvement of $32.7\%$ in inference time.
Limitations: (1) Mjölnir is not effective in attacking perturbations that are not based on gradient diffusion (noise perturbation) such as representation perturbation. (2) The overall effectiveness of Mjölnir in reconstructing privacy datasets is also bounded by the selected subsequent Gradient Leakage Attacks. (3) Mjölnir is used to attack CNN and (or) DNN-based models. Experimental results suggest that it struggles with transformer-based models due to their complex structure and larger number of parameters, making it difficult to accurately denoise gradients and recover data.
@@
Defense Strategies:Possible defense strategies against Mjölnir can be approached by preserving the privacy of original data, preserving the target attack models, and shared gradients protection. Mjölnir effectively leakages privacy data by attacking the perturbed shared gradients, which suggests the possible defense approaches against Mjölnir should either be non-perturbation gradient protection methods or non-gradient privacy-preserving methods. For instance, processing of the participant’s local training model in depth and layers to weaken the performance of Gradient Leakage Attacks, encryption on original local privacy data before training, and various cryptography-based protections on shared gradients (Huang et al. 2021; Zhang et al. 2022; Rodríguez-Barroso et al. 2023). Cryptography-based approaches always cause more computation time and storage (Zhang et al. 2020), thus lightweight defense strategies that can serve as substitutes for invalid perturbation protection still need future explorations.
Conclusion